OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the line is undefined, state so. (9, -6) and (5, 6)

OpenStudy (anonymous):

6 +6 ----- = -3 -4

OpenStudy (anonymous):

nope correct answer is 60

OpenStudy (anonymous):

(Y2-Y1)/(x2-x1)= (6-(-6))/(5-9)=(12/(-4))=(-3)

OpenStudy (anonymous):

poliboy u r insane :)

OpenStudy (anonymous):

poliboy imagine a line with slope 60 haha lol rofl :D

OpenStudy (anonymous):

sometimes maths is so funny lol :D

OpenStudy (anonymous):

don't look at me it is what the book gave for the answer

OpenStudy (anonymous):

burn it

OpenStudy (anonymous):

books can be wrong

OpenStudy (anonymous):

its lying to u